OpenStudy (anonymous):

what is the focus of the parabola given by the equation y= -1/28 x^2 ?

OpenStudy (anonymous):

please explain, thankyouuu!

OpenStudy (anonymous):

uhh. not really. this whole lesson confuses me! I know what a vertex is, I just dont know what it is in this problem

OpenStudy (anonymous):

but wait how do I find the focus?

OpenStudy (anonymous):

http://www.math-prof.com/Alg2/Alg2_Ch_40.asp

OpenStudy (anonymous):

so I try and get the x^2 by itself?

OpenStudy (anonymous):

\[ y= -\frac{1}{28}x^2 \] \[-28y=x^2\] \[4p=-28\] \[p=-7\]

OpenStudy (saifoo.khan):

Hey! How you did the 3rd step? @satellite73

OpenStudy (saifoo.khan):

Nice question @mother123

OpenStudy (anonymous):

wait yeah how did you do the third step, Im lost.

OpenStudy (saifoo.khan):

When Satellite's near, we don't need to worry! :D

OpenStudy (anonymous):

haha so the answer is (-7,0) or (0,-7) Those are the choices I have for my multiple choice

OpenStudy (anonymous):

your goal is to make it look like \(4py=x^2\) so you can find \(p\) and then find the focus

OpenStudy (anonymous):

once we have \(-28y=x^2\) we know \(4p=-28\) and so we know \(p=-7\) but we are not done yet

OpenStudy (saifoo.khan):

so 4py=x^2 is something like a formula?

OpenStudy (anonymous):

yes old school formula or more generally \(4p(y-k)=(x-h)^2\)

OpenStudy (anonymous):

now to actually find the focus, we draw a picture 

OpenStudy (anonymous):

since the parabola has vertex at origin and opens down, we see the focus is \((0,-7)\)

OpenStudy (anonymous):

so then the focus is (0,-7)?

OpenStudy (saifoo.khan):

Perfect!!! Last question: Now what's the directix? @satellite73

OpenStudy (anonymous):

ohh I get it a little more! wait I have another one, can you tell me If Im doing it right?

OpenStudy (saifoo.khan):

I will try that too!

OpenStudy (anonymous):

evidently since the directrix is a line same distance from the vertex as the focus, it should be \(y=7\)

OpenStudy (anonymous):

x=3(y-1)^2+2

OpenStudy (anonymous):

do I divide everything by 3?

OpenStudy (saifoo.khan):

What's the exact meaning of directix?

OpenStudy (saifoo.khan):

Why is it there?

OpenStudy (saifoo.khan):

@mother123 , you should have x^2 as the subject if in not wrong?

OpenStudy (saifoo.khan):

im*

OpenStudy (anonymous):

one definition of the parabola is all points whose distance from a fixed point (focus) and a line (directrix) are equal

OpenStudy (saifoo.khan):

Satellite is the superman here: http://femour.com/wp-content/uploads/2012/07/tumblr_m7h2z1Basv1qiekquo1_500.jpg

OpenStudy (anonymous):

idk. there is no x^2 in the problem

OpenStudy (anonymous):

\[x=3(y-1)^2+2\] \[x-2=3(y-1)^2\] \[\frac{1}{3}(x-2)=(y-1)^2\]

OpenStudy (saifoo.khan):

What's the original question?

OpenStudy (anonymous):

don't lose track of the goal you want to make it look like \(4p(x-h)=(y-k)^2\) or \[4p(y-k)=(x-h)^2\]

OpenStudy (anonymous):

parabola that opens up or down, with vertex at \((h,k)\) will look like \[4p(y-k)=(x-h)^2\] and parabola that opens left or right with vertex \((h,k)\) will look like \[4p(x-h)=(y-k)^2\]

OpenStudy (anonymous):

how do you get 1/3 to be a 4?

OpenStudy (saifoo.khan):

i see..

OpenStudy (anonymous):

the (h\) always goes with the \(x\) and the \(k\) always goes with the \(y\) make it iin to this form to get what you want

OpenStudy (anonymous):

@mother123 you do not make \(\frac{1}{3}=4\) (which would be a miracle) you make \(4p=\frac{1}{3}\) and solve for \(p\)

OpenStudy (saifoo.khan):

so p = 1/12

OpenStudy (anonymous):

This is the hardest thing I've ever done. Im so confused.

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

idk if this is right?